OpenStudy (anonymous):

how do you figure out during what year it will double if the compounding is continous with the given info: invest 10,000 dollars at 4.5% interest. What is in your account after 20 years

OpenStudy (nottim):

So, Compound interest. You know the formula?

OpenStudy (anonymous):

10,000 * e^(0.045*20)

OpenStudy (anonymous):

Use a calculator of some kind that has the capability to compute exponents of e.

OpenStudy (anonymous):

yes i know how to find the continous compound but i dont understnad how to find out during what year it will double if the compounding is continous

OpenStudy (anonymous):

That means you need to set the equation to 2 times original value and find "t". 10,000 * e^(0.045*t) = 20,000 e^(0.045*t) = 2 0.045*t*ln(e) = ln(2) 0.045*t = ln(2) t = ln(2)/0.045 t = 15.4 So, the principal will double in the 16th year.